A force produce an acceleration of 16m/s2 body of mass 0.5kg and an accleration of 4.0m/s2 in another body.If both  the bodies are fastended together then how much acceleration will be produced by this force?

Grade:12th pass

1 Answers

Shailendra Kumar Sharma
6 years ago
F= maFor 0.5 a= 16m/S2I.e. F= 32NAnother body a= 4m/S2So m= 8Now when tied togetherm= 8.5So a= 32/8.5=3.76m/s2